What is the linear equation containing the point (-4, -1) and perpendicular to the line through (1, 4) and (3, 5) in general form?
Answer by MathDazed(34)

First we find the slope of the perpendicular line:
Using (1,4) and (3,5)
slope = (y[2]-y[1])/(x[2]-x[1])
= (5 - 4)/(3 - 1)
= 1/2
If the line is perpendicular to the above then it's slope is -2 (negative reciprocal)
Using the slope intercept formula, you have y = mx + b where m is the slope and b is the y-intercept
y = -2x + b
We will now use the point (-4,-1)
-1 = -2(-4) + b
-1 = 8 + b
-1 - 8 = b
b = -9
Therefore, the linear equation is y = -2x -9 or 2x + y = -9